# Scalable Estimation of Multinomial Response Models

### v4

@@ 1 Introduction
In general, this paper contributes to the expanding literature on high-dimensional demand estimation in statistics and marketing: (Braun and McAuliffe, 2010; Chiong and Shum, 2019; Smith and Allenby, 2019; Loaiza-Maya and Nibbering, 2022; Jiang et al., 2024; Iaria and Wang, 2024; Ershov et al., 2024; Amano et al., 2018). To incorporate latent consideration sets, it is necessary to generalize the standard multinomial response model by conditioning the distribution of responses on a latent subject-specific consideration set, which is drawn from the power set of $\{1,2,\ldots,J\}$. This results in a mixture of multinomial models based on a probability distribution over consideration sets. However, the exponential size of this power set renders the estimation of this mixture of multinomial response models computationally infeasible in general. Moreover, the proposed method can be interpreted as a generalized multinomial logit (MNL) model, with “structural zeros” incorporated in the first layer of its hierarchical structure. In the field of biostatistics, methodologies have been extensively explored to estimate microbial compositions that account for the sparsity due to excessive zero counts (e.g. Aitchison, 1982; Martín-Fernández et al., 2015; Liu et al., 2020; Cao et al., 2020; Paulson et al., 2013; Chen and Li, 2016; Tang and Chen, 2019). More recently, Zeng et al. (2023) introduced a zero-inflated probabilistic PCA model designed for high-dimensional, sparse microbiome data sets. Although our paper focuses on a different problem, the proposed method has the potential to be applied in similar contexts, as we discuss in the concluding section.
@@ 2.1 The latent consideration sets
To fix notation, let $\mathcal{C}$ represent the collection of all possible consideration sets, which corresponds to the power set of $\mathcal{J}=\{1,\ldots,J\}$, excluding the empty set. The consideration set for subject $i$ is indicated by $\mathcal{C}_{i}=c$, where $c\in\mathcal{C}$. For example, when $J=3$, $\mathcal{C}=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\text{and}\,\{1,2,3\}\}$, and $c$ is one of these elements. Furthermore, by $\bm{C}_{i}=(C_{i1},\ldots,C_{iJ})^{\prime}$, we mean a $J\times 1$ multivariate binary vector where $C_{ij}=1$ if category $j$ is in the consideration set, and 00 otherwise. In the example of $J=3$, $\mathcal{C}_{i}=\{1\}$ is equivalent to $\bm{C}_{i}=(1,0,0)^{\prime}$ and $\mathcal{C}_{i}=\{1,3\}$ is equivalent to $\bm{C}_{i}=(1,0,1)^{\prime}$ etc. In the following, we use the two notations interchangeably depending on the context. Researchers sometimes include an outside option in the model that is always considered by each subject. We can incorporate this into our framework by adding a $(J+1)$th category and fixing $C_{iJ+1}=1$ for all $i$. Our goal is to put a probability distribution on $\mathcal{C}$ that is rich enough to accommodate dependencies while maintaining scalability.
@@ 3 Theoretical results
[…]
@@
We now address the possibility that multiple parameter pairs $(\bm{\theta},\bm{\pi})$ may be consistent with the true response probabilities. This relates to the issue of partial identification (Masatlioglu et al., 2012, Cattaneo et al., 2020, Barseghyan et al., 2021a, Lu, 2022), where point identification holds only under specific conditions (Dardanoni et al., 2020, Abaluck and Adams-Prassl, 2021, Barseghyan et al., 2021b). Following Aguiar and Kashaev (2024), we impose the assumption that the panel is sufficiently long and that random effects are absent. Under these conditions, we show that the two sources of variation in responses—differences in utility and differences in consideration sets—can be separately identified. Formally, we show in the next theorem that the posterior distribution contracts to within an arbitrarily small ball around $(\bm{\beta}^{*},\bm{\pi}^{*})$ under the distance function $d((\bm{\beta},\bm{\pi}),(\bm{\beta}^{\prime},\bm{\pi}^{\prime}))=\max\{||\bm{\pi}-\bm{\pi}^{\prime}||_{1},||\bm{\beta}-\bm{\beta}^{\prime}||_{2}\}$.
@@ 4.3 Numerical illustration
We illustrate posterior probabilities of consideration sets on synthetic panel data with $n=100$ subjects observed over $T\in\{1,2,\ldots,15\}$ time periods. We let $J=4$ and give the $152^{J}-1=$ 15 consideration sets in the first column of Table 1. In the table we report the posterior probabilities of each possible consideration set for a randomly chosen subject $i$ whose true consideration set is $\mathcal{C}_{i}^{*}=\{1,3,4\}$.
@@ 5 Monte Carlo Simulation
We demonstrate the sampling performance of the proposed approach through simulation studies containing $J=4$ and $J=101$ alternatives. The goal is to empirically validate the findings of Theorem 2 and demonstrate that the proposed approach can effectively assess consideration dependence. In the Supplementary Material, we conduct additional experiments under autocorrelated covariates, random effects, and time-varying true consideration sets. In general, the experiments show posterior consistency in the estimation of $\bm{\theta}=(\bm{\beta},\bm{D})$ and $\bm{\pi}$, and that the restrictive approach with $K=1$ produces larger mean-square-errors and biases.
@@ 5.1 $J=4$
We let $T_{i}=T$ for all $i$. In one case we set $T=5$ and in the other $T=15$. The latter satisfies the length condition of Theorem 2. In simulating the data, we first specify the distribution of the consideration sets $\bm{\pi}^{*}=\{\pi^{*}_{c}=\Pr(\mathcal{C}_{i}=c):c\in\mathcal{C}\}$. We induce dependence in product consideration by letting the first two and last two products have a relatively high probability of being considered together: $\pi^{*}_{\{1,2\}}=\pi^{*}_{\{3,4\}}=0.25$. As motivation, the first two products might represent non-vegetarian options, and the last two vegetarian. The other 13 consideration sets $c\in\mathcal{C}$ are given a probability of 0.0385. Figure 1 shows $\bm{\pi}^{*}$ in red. Given this $\bm{\pi}^{*}$, we generate the true consideration sets $\mathcal{C}_{i}^{*}$, for $i=1,\ldots,n$. We then generate outcomes from the logit model with $V_{ijt}=\delta^{*}_{j}+\beta^{*}x_{ijt},$ letting $(\delta_{1}^{*},\delta_{2}^{*},\delta_{3}^{*},\delta_{4}^{*})^{\prime}=(1.0,0.5,-1.0,0)^{\prime}$ and $\beta^{*}=1$, and $x_{ijt}\overset{iid}{\sim}N(0,1)$.
We compare the performance between the proposed infinite mixture of independent consideration models ($K=\infty$) and the model that assumes independent consideration ($K=1$) over 50 replicated data sets. The results are given in Table 2 where we report the mean-square-error (MSE), posterior standard deviation, and empirical coverage of the equal-tailed 95% credible intervals and the computational time. As $n$ increases, the posterior of the response parameter $\bm{\beta}=(\delta_{1},\delta_{2},\delta_{3},\beta)$ and the distribution of the consideration sets $\bm{\pi}$ is close to the true values, even when $T=5$. In contrast, when $K=1$, we do not observe sufficient evidence of posterior consistency. In particular, MSEs are an order of magnitude larger in some cases than those under $K=\infty$, due to misspecification. When $T$ increases to $15$, the MSEs/SDs become smaller for both $K=\infty$ and $K=1$, but there are distortions in the coverage for $K=1$.
@@
$(\infty,5)$ $50$ 0.029 ( 0.16 ; 0.92 ) 0.28 ( 0.42 ; 0.90 ) 0.299 ( 0.43 ; 0.90 ) 0.098 ( 0.34 ; 0.96 ) 0.46 ( 0.03 ) 1.68
$200$ 0.006 ( 0.08 ; 0.96 ) 0.047 ( 0.22 ; 0.98 ) 0.041 ( 0.22 ; 0.98 ) 0.034 ( 0.17 ; 0.94 ) 0.266 ( 0.01 ) 5.7
$(\infty,15)$ $50$ 0.007 ( 0.09 ; 0.94 ) 0.033 ( 0.2 ; 0.98 ) 0.037 ( 0.2 ; 0.94 ) 0.028 ( 0.18 ; 0.98 ) 0.36 ( 0.01 ) 2.19
$200$ 0.002 ( 0.04 ; 0.92 ) 0.011 ( 0.1 ; 0.92 ) 0.01 ( 0.1 ; 0.92 ) 0.007 ( 0.09 ; 0.92 ) 0.191 ( 0.01 ) 7.54
$(1,5)$ $50$ 0.024 ( 0.15 ; 0.96 ) 0.787 ( 0.4 ; 0.46 ) 0.927 ( 0.4 ; 0.42 ) 0.094 ( 0.33 ; 0.94 ) 0.728 ( 0.03 ) 1.4
$200$ 0.01 ( 0.07 ; 0.8 ) 0.894 ( 0.19 ; 0.00 ) 0.949 ( 0.2 ; 0.02 ) 0.032 ( 0.17 ; 0.94 ) 0.721 ( 0.01 ) 4.63
$(1,15)$ $50$ 0.007 ( 0.09 ; 0.94 ) 0.068 ( 0.21 ; 0.86 ) 0.073 ( 0.21 ; 0.90 ) 0.032 ( 0.18 ; 0.98 ) 0.402 ( 0.01 ) 1.98
$200$ 0.002 ( 0.04 ; 0.94 ) 0.029 ( 0.11 ; 0.76 ) 0.023 ( 0.11 ; 0.78 ) 0.011 ( 0.09 ; 0.88 ) 0.261 ( 0.01 ) 6.17
  • For $\beta$ and $\delta$, we show the average of the mean squared errors (MSEs), using the posterior means as point estimators. In parentheses, the average posterior standard deviations (SDs) and empirical coverages of 95% credible intervals are shown. For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ with the posterior SDs in the parentheses. Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM.
@@ 5.2 $J=100$
We now consider a high-dimensional scenario with $J=100$. One mechanism by which the dependence of consideration among categories can be induced is through multiple latent subpopulations of subjects having different probabilities of consideration. Within a subpopulation, considerations are independent across categories. However, marginalizing out the latent subpopulation indicator, one obtains dependence in those category considerations. We generate the data with two subpopulations. To generate the true consideration set of a given subject, we used a Bernoulli distribution with attention probability 0.05 for each category except for categories 10, 30, 50, 70, and 90 for the first subpopulation ($i=1,\ldots,50$) where the Bernoulli attention probability was set to 0.8. For the remaining subjects in the second subpopulation ($i=51,\ldots,100$), the Bernoulli probability was set at 0.05 except for categories 20, 40, 60, 80, and 100 where the probability was set to 0.8. Conditional on the true consideration sets, we generated the responses as in the case with $J=4$.
(a) $T=1$
(b) $T=10$
(c) $T=200$
Figure 2: Upper panels: Posterior mean of $1\{j\in\mathcal{C}_{1}\}$ for each item $j=1,\ldots,J$. Filled square(circle) if the category $j$ within the true consideration set $\mathcal{C}_{1}^{*}=\{10,30,50,70,75,88,90\}$ has(not) been chosen by $T$. Unfilled circle if it is outside $\mathcal{C}_{1}^{*}$. Lower panels: similarity matrices. The results are based on the simulated data with $J=100$, $T=\{1,10,200\}$, and $n=100$. The time (seconds) taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM are 53.5, 97.1, and 1788.0, respectively.
Because in this case there are $12^{100}-$ 1 support points in the distribution of the consideration sets, it is not possible to show the entire distribution as in the case of $J=4$ before. We focus on the estimation results regarding subject 1 whose true consideration set contains 7 categories: $\mathcal{C}_{1}^{*}=\{10,30,50,70,75,88,90\}$. The upper panels of Figure 2 show the posterior probabilities that $\mathcal{C}_{1}$ include each of the categories (filled square/circle if the particular category in the true consideration set has/has not been chosen by $T$; unfilled circle if it is outside the true consideration set). The lower panels of Figure 2 are the $n\times n$ “similarity matrices”. These give the posterior probability that a given subject in a particular row $k$ is in the same cluster as another subject at a specific column $i$ which is computed as the posterior probability of the event $\{S_{k}=S_{i}\}$. This probability ranges from zero (light blue) to one (dark blue).
We again observe posterior concentration toward the true consideration set as $T$ increases. Even with a relatively short $T=10$, category 70, which subject 1 has not yet chosen, has a relatively high posterior probability of being an element of the subject’s consideration set. This occurs because subject 1 is clustered with others whose observed choices include category 70. By $T=10$, the estimated similarity matrix closely aligns with the true clustering structure, leading to consistent consideration set estimates within each cluster. For $T=200$, a span that approximately satisfies the identification condition in Theorem 2, the clustering structure is recovered with high accuracy.

### v5

@@ 1 Introduction
In general, this paper contributes to the expanding literature on high-dimensional demand estimation in statistics and marketing: (Braun and McAuliffe, 2010; Chiong and Shum, 2019; Smith and Allenby, 2019; Loaiza-Maya and Nibbering, 2022; Jiang et al., 2024; Iaria and Wang, 2024; Ershov et al., 2024; Amano et al., 2018). To incorporate latent consideration sets, it is necessary to generalize the standard multinomial response model by conditioning the distribution of responses on a latent subject-specific consideration set, which is drawn from the power set of $\{1,2,\ldots,J\}$. This results in a mixture of multinomial models based on a probability distribution over consideration sets. However, the exponential size of this power set renders the estimation of this mixture of multinomial response models computationally infeasible in general. Moreover, the proposed method can be interpreted as a generalized multinomial logit (MNL) model, with “structural zeros” incorporated in the first layer of its hierarchical structure. In the field of biostatistics, methodologies have been extensively explored to estimate microbial compositions that account for the sparsity due to excessive zero counts (e.g. Aitchison, 1982; Mart´ın-Fernández et al., 2015; Liu et al., 2020; Cao et al., 2020; Paulson et al., 2013; Chen and Li, 2016; Tang and Chen, 2019). More recently, Zeng et al. (2023) introduced a zero-inflated probabilistic PCA model designed for high-dimensional, sparse microbiome data sets. Although our paper focuses on a different problem, the proposed method has the potential to be applied in similar contexts, as we discuss in the concluding section.
@@ 2.1 The latent consideration sets
To fix notation, let $\mathcal{C}$ represent the collection of all possible consideration sets, which corresponds to the power set of $\mathcal{J}=\{1,\ldots,J\}$, excluding the empty set. The consideration set for subject $i$ is indicated by $\mathcal{C}_{i}=c$, where $c\in\mathcal{C}$. For example, when $J=3$, $\mathcal{C}=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\text{and}\,\{1,2,3\}\}$, and $c$ is one of these elements. Furthermore, by $\bm{C}_{i}=(C_{i1},\ldots,C_{iJ})^{\prime}$, we mean a $J\times 1$ multivariate binary vector where $C_{ij}=1$ if category $j$ is in the consideration set, and 0 otherwise. In the example of $J=3$, $\mathcal{C}_{i}=\{1\}$ is equivalent to $\bm{C}_{i}=(1,0,0)^{\prime}$ and $\mathcal{C}_{i}=\{1,3\}$ is equivalent to $\bm{C}_{i}=(1,0,1)^{\prime}$ etc. In the following, we use the two notations interchangeably depending on the context. Researchers sometimes include an outside option in the model that is always considered by each subject. We can incorporate this into our framework by adding a $(J+1)$th category and fixing $C_{iJ+1}=1$ for all $i$. Our goal is to put a probability distribution on $\mathcal{C}$ that is rich enough to accommodate dependencies while maintaining scalability.
@@ 3 Theoretical results
Given a $K\in\mathbb{Z}^{+}$, define $\bm{\phi}_{1:K}=\{\omega_{h},\bm{q}_{h}:h=1,\ldots,K\}$, the collection of all component-specific parameters, where $\bm{q}_{h}=(q_{h1},\ldots,q_{hJ})^{\prime}$. Note that by Lemma 1, there exist $\{K,\tilde{\bm{\phi}}_{1:K}\}$, which may not be unique, such that $\pi_{c}^{*}=\sum_{h=1}^{K}\tilde{\omega}_{h}\left\{\prod_{j\in c}\tilde{q}_{hj}\prod_{j\notin c}\big{(}1-\tilde{q}_{hj}\big{)}\right\},$  $\text{ for all }c\in\mathcal{C},$ and the KL divergence is zero at $\{\bm{\theta}^{*},K,\tilde{\bm{\phi}}_{1:K}\}$. In the following lemma, we establish that the KL divergence can be made arbitrarily small in sufficiently small neighborhoods of $(\bm{\theta}^{*},\tilde{\bm{\phi}}_{1:K})$. Define the model induced probability for a consideration set $c\in\mathcal{C}$: $\pi(c|K,\bm{\phi}_{1:K})=\sum_{h=1}^{K}\omega_{h}\prod_{j\in c}q_{hj}\prod_{j\notin c}(1-q_{hj}),$ and the model induced marginal response probability as
@@
[…]
@@ 4.3 Numerical illustration
We illustrate posterior probabilities of consideration sets on synthetic panel data with $n=100$ subjects observed over $T\in\{1,2,\ldots,15\}$ time periods. We let $J=4$ and give the $2^{J}-1=15$ consideration sets in the first column of Table 1. In the table we report the posterior probabilities of each possible consideration set for a randomly chosen subject $i$ whose true consideration set is $\mathcal{C}_{i}^{*}=\{1,3,4\}$.
@@ 5 Monte Carlo Simulation
We demonstrate the sampling performance of the proposed approach through simulation studies first with $J=4$ alternatives, where it is possible to enumerate all the support points in $\bm{\pi}$, and then extend the study to a high-dimensional case with $J=100$. The goal is to empirically validate the findings of Theorem 2 and demonstrate that the proposed approach can effectively assess consideration dependence. In the Supplementary Material, we conduct additional experiments under autocorrelated covariates, random effects, and time-varying true consideration sets. In general, the experiments show posterior consistency in the estimation of $\bm{\theta}=(\bm{\beta},\bm{D})$ and $\bm{\pi}$, and that the restrictive approach with $K=1$ produces larger root mean squared errors and biases.
@@ 5.1 $J=4$
We let $T_{i}=T$ for all $i$. In one case we set $T=5$ and in the other $T=15$. The latter satisfies the length condition of Theorem 2. In simulating the data, we first specify the distribution of the consideration sets $\bm{\pi}^{*}=\{\pi^{*}_{c}=\Pr(\mathcal{C}_{i}=c):c\in\mathcal{C}\}$. We induce dependence in product consideration by letting the first two and last two products have a relatively high probability of being considered together: $\pi^{*}_{\{1,2\}}=\pi^{*}_{\{3,4\}}=0.25$. As motivation, the first two products might represent non-vegetarian options, and the last two vegetarian. The other 13 consideration sets $c\in\mathcal{C}$ are given a probability of 0.0385. Figure 1 shows $\bm{\pi}^{*}$ in red. Given this $\bm{\pi}^{*}$, we generate the true consideration sets $\mathcal{C}_{i}^{*}$, for $i=1,\ldots,n$. We then generate outcomes from the logit model with $V_{ijt}=\delta^{*}_{j}+\beta^{*}x_{ijt},$ letting $(\delta_{1}^{*},\delta_{2}^{*},\delta_{3}^{*},\delta_{4}^{*})^{\prime}=(1.0,0.5,-1.0,0)^{\prime}$ and $\beta^{*}=1$, and $x_{ijt}\overset{iid}{\sim}N(0,1)$. We let $n\in\{50,100\}$.
We compare the performance between the proposed infinite mixture of independent consideration models ($K=\infty$) and the model that assumes independent consideration ($K=1$) over 200 replicated data sets. The results are given in Table 2 where we report the root mean squared error (RMSE) for the response parameter $\bm{\beta}=(\delta_{1},\delta_{2},\delta_{3},\beta)$ as well as the $L_{1}$ norm between the posterior mean and the truth for the distribution of the consideration sets $\bm{\pi}$ (L1-error), their Monte Carlo errors (MCE), the posterior standard deviation (SD), the empirical standard deviation (ESD), the empirical coverage of the equal-tailed 95% credible intervals (Cov), and the computational time. The MCE quantifies the precision for the performance criterion. The MCEs are negligible, allowing for valid comparisons based on the $200$ replications. As $n$ increases, the posterior of $\bm{\beta}$ and $\bm{\pi}$ contracts to the true values, even when $T=5$, indicated by the smaller RMSEs and L1-errors as well as SDs. In contrast, when $K=1$, we do not observe sufficient evidence of posterior consistency. The RMSEs and L1-error are much larger in some cases than those under $K=\infty$, due to misspecification. When $T$ increases to $15$, a value that satisfies the identifying condition in Theorem 2, the RMSEs/L1-errors/SDs become smaller for both $K=\infty$ and $K=1$, but for $K=1$, they are larger, and there are distortions in the coverage. Finally, our approach ($K=\infty$) delivers good coverages in general. The SDs are similar to ESDs, indicating that the posterior standard deviations provide a good representation of the sampling variability of the posterior means.
@@
RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov L1-error (MCE) SD (ESD)       Cov
$(\infty,5)$ $50$ 0.168 ( 0.01 ) 0.16 ( 0.167 ) 0.94 0.464 ( 0.024 ) 0.44 ( 0.442 ) 0.94 0.47 ( 0.023 ) 0.44 ( 0.442 ) 0.94 0.328 ( 0.018 ) 0.34 ( 0.329 ) 0.96 0.446 ( 0.01 ) 0.03 ( 0.028 ) 0.94 1.84
$100$ 0.116 ( 0.005 ) 0.11 ( 0.115 ) 0.98 0.361 ( 0.021 ) 0.31 ( 0.32 ) 0.91 0.359 ( 0.021 ) 0.31 ( 0.324 ) 0.92 0.235 ( 0.01 ) 0.25 ( 0.227 ) 0.96 0.366 ( 0.006 ) 0.02 ( 0.022 ) 0.96 3.37
$(\infty,15)$ $50$ 0.092 ( 0.004 ) 0.09 ( 0.093 ) 0.93 0.194 ( 0.01 ) 0.21 ( 0.189 ) 0.96 0.185 ( 0.01 ) 0.2 ( 0.179 ) 0.95 0.171 ( 0.009 ) 0.17 ( 0.172 ) 0.96 0.358 ( 0.005 ) 0.03 ( 0.023 ) 0.96 2.51
$100$ 0.062 ( 0.003 ) 0.06 ( 0.062 ) 0.97 0.132 ( 0.007 ) 0.14 ( 0.13 ) 0.96 0.136 ( 0.007 ) 0.14 ( 0.136 ) 0.97 0.117 ( 0.006 ) 0.12 ( 0.116 ) 0.97 0.294 ( 0.003 ) 0.02 ( 0.018 ) 0.93 4.84
$(1,5)$ $50$ 0.161 ( 0.008 ) 0.15 ( 0.158 ) 0.94 0.883 ( 0.028 ) 0.4 ( 0.442 ) 0.49 0.943 ( 0.029 ) 0.4 ( 0.473 ) 0.47 0.33 ( 0.019 ) 0.33 ( 0.331 ) 0.96 0.847 ( 0.006 ) 0.03 ( 0.024 ) 0.5 1.62
$100$ 0.112 ( 0.005 ) 0.11 ( 0.104 ) 0.91 0.949 ( 0.024 ) 0.28 ( 0.336 ) 0.18 0.981 ( 0.027 ) 0.28 ( 0.355 ) 0.18 0.243 ( 0.012 ) 0.24 ( 0.239 ) 0.94 0.876 ( 0.007 ) 0.02 ( 0.02 ) 0.17 2.81
$(1,15)$ $50$ 0.092 ( 0.004 ) 0.09 ( 0.092 ) 0.93 0.283 ( 0.016 ) 0.22 ( 0.23 ) 0.92 0.272 ( 0.017 ) 0.22 ( 0.218 ) 0.92 0.182 ( 0.01 ) 0.18 ( 0.178 ) 0.94 0.714 ( 0.002 ) 0.02 ( 0.018 ) 0.89 2.24
$100$ 0.062 ( 0.003 ) 0.06 ( 0.062 ) 0.96 0.206 ( 0.01 ) 0.15 ( 0.151 ) 0.87 0.199 ( 0.012 ) 0.15 ( 0.159 ) 0.85 0.133 ( 0.007 ) 0.12 ( 0.122 ) 0.95 0.706 ( 0.001 ) 0.01 ( 0.013 ) 0.82 4.24
  • For $\beta$ and $\delta$, for each case, we show the estimated root mean squared error (RMSE), using the posterior means as point estimator. In parenthesis, the jackknife estimate of Monte Carlo Error (MCE) for the RMSE is presented. Next, the average of the posterior standard deviations (SD) is shown with the empirical standard deviation (ESD) of the posterior mean in the parenthesis. Third, the empirical coverage (Cov) of 95% credible interval is given.
  • Letting $\hat{\theta}_{r}$ denote the posterior mean from replication $r$ $(r=1,\ldots,R)$ and letting $\theta^{\ast}$ denote the true value, we summarize finite-sample accuracy and variability using:
    • $RMSE=\sqrt{\frac{1}{R}\sum_{r=1}^{R}(\hat{\theta}_{r}-\theta^{*})^{2}}$,
    • $MCE(\widehat{RMSE})=\sqrt{\frac{R-1}{R}\sum_{r=1}^{R}(RMSE_{(-r)}-\overline{RMSE_{(-r)}})^{2}}$, where $RMSE_{(-r)}$ is the RMSE estimated with the $r$th replicate removed and $\overline{RMSE_{(-r)}}=\frac{1}{R}\sum_{r=1}^{R}RMSE_{(-r)}$,
    • $ESD=\sqrt{\frac{1}{R-1}\sum_{r=1}^{R}(\hat{\theta}_{r}-\bar{\hat{\theta}})^{2}}$, where $\bar{\hat{\theta}}=\frac{1}{R}\sum_{r=1}^{R}\hat{\theta}_{r}$.
  • For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ (L1-error). In the parenthesis, we show its jackknife estimate of MCE. The SDs, ESDs, and Covs are averaged over the 15 elements in $\pi$.
  • Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM. The study is based on $R=200$ replications. 2,000 MCMC draws are obtained for each replication. The average of the inefficiency factors is around 6.6 with standard deviation 1.2.
@@ 5.2 $J=100$
We now consider a high-dimensional scenario with $J=100$ alternatives. One mechanism by which the dependence of consideration among categories can be induced is through multiple latent subpopulations of subjects having different probabilities of consideration. Within a subpopulation, considerations are independent across categories. However, marginalizing out the latent subpopulation indicator, one obtains dependence in those category considerations. We generate the data with two subpopulations. To generate the true consideration set of a given subject, we used a Bernoulli distribution with attention probability 0.05 for each category except for categories 10, 30, 50, 70, and 90 for the first subpopulation ($i=1,\ldots,n/2$) where the attention probability was set to 0.8. For the remaining subjects in the second subpopulation ($i=n/2+1,\ldots,n$), the Bernoulli probability was set at 0.05 except for categories 20, 40, 60, 80, and 100 where the probability was set to 0.8. Conditional on the true consideration sets, we generated the responses as in the case with $J=4$ with $\delta_{j}^{*}=0,\ j=1,\ldots,J-1$.
Because in this case there are $2^{100}-1$ support points in $\bm{\pi}$, it is not possible to show the entire distribution as in the case of $J=4$. Also, there are 99 $\delta_{j}$’s to estimate. Hence, in Table 3, we report the results only for the slope $\beta$ as well as $\delta_{97},\delta_{98},$ and $\delta_{99}$. The results are based on 200 replications. The general observations from the small $J$ simulation still hold: as $n$ increases, the RMSEs/SDs of $\bm{\beta}$ become smaller even when $T=5$, supporting posterior consistency. For $T=200$, a span that approximately satisfies the identification condition in Theorem 2, our approach also results in good coverages.
Table 3: Simulation results with $J=100$
$(K,T)$ $n$ $\beta$ $\delta_{97}$ $\delta_{98}$ $\delta_{99}$ Time
RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov
$(\infty,5)$ $50$ 0.101 ( 0.004 ) 0.08 ( 0.083 ) 0.85 0.82 ( 0.036 ) 0.94 ( 0.803 ) 0.95 0.811 ( 0.049 ) 0.91 ( 0.783 ) 0.95 0.85 ( 0.048 ) 0.93 ( 0.837 ) 0.90 3.32
$100$ 0.09 ( 0.004 ) 0.06 ( 0.057 ) 0.79 0.777 ( 0.037 ) 0.79 ( 0.741 ) 0.93 0.793 ( 0.037 ) 0.79 ( 0.765 ) 0.93 0.816 ( 0.04 ) 0.81 ( 0.8 ) 0.92 7.29
$(\infty,200)$ $50$ 0.014 ( 0.001 ) 0.01 ( 0.014 ) 0.92 0.261 ( 0.014 ) 0.29 ( 0.239 ) 0.84 0.232 ( 0.01 ) 0.25 ( 0.211 ) 0.86 0.231 ( 0.014 ) 0.35 ( 0.218 ) 0.90 216.4
$100$ 0.01 ( 0.001 ) 0.01 ( 0.009 ) 0.97 0.145 ( 0.01 ) 0.14 ( 0.128 ) 0.93 0.127 ( 0.006 ) 0.13 ( 0.11 ) 0.96 0.165 ( 0.011 ) 0.14 ( 0.153 ) 0.92 417.9
  • For $\beta$ and $\delta$, for each case, we show the estimated root mean squared error (RMSE), using the posterior means as point estimator. In parenthesis, the jackknife estimate of Monte Carlo Error (MCE) for the RMSE is presented. Next, the average of the posterior standard deviations (SD) is shown with the empirical standard deviation (ESD) of the posterior mean in the parenthesis. Third, the empirical coverage (Cov) of 95% credible interval is given. Time is the average minutes taken for sampling 1,000 MCMC draws. The study is based on $R=200$ replications. 3,000 MCMC draws are obtained for each replication. The average of the inefficiency factors is around 3.26 with standard deviation 0.79.